# https://humanaigc.github.io/emote-portrait-alive/ EMO: Emote Portrait Alive - Generating Expressive Portrait Videos with Audio2Video Diffusion Model under Weak Conditions

### v1

@@ Abstract
In this work, we tackle the challenge of enhancing the realism and expressiveness in talking head video generation by focusing on the dynamic and nuanced relationship between audio cues and facial movements. We identify the limitations of traditional techniques that often fail to capture the full spectrum of human expressions and the uniqueness of individual facial styles. To address these issues, we propose EMO, a novel framework that utilizes a direct audio-to-video synthesis approach, bypassing the need for intermediate 3D models or facial landmarks. Our method ensures seamless frame transitions and consistent identity preservation throughout the video, resulting in highly expressive and lifelike animations. Experimental results demonsrate that EMO is able to produce not only convincing speaking videos but also singing videos in various styles, significantly outperforming existing state-of-the-art methodologies in terms of expressiveness and realism.
@@ 1 Introduction
In recent years, the field of image generation has witnessed remarkable advancements, largely attributed to the emergence and success of Diffusion Models [8, 20, 25, 4, 16]. These models, celebrated for their ability to produce high-quality images, owe their prowess to extensive training on large-scale image datasets and a progressive generation approach. This innovative methodology enables the creation of images with unparalleled detail and realism, setting new benchmarks in the domain of generative models. The application of Diffusion Models has not been confined to still images alone. A burgeoning interest in video generation has led to the exploration of these models’ potential in crafting dynamic and compelling visual narratives [6, 9]. These pioneering efforts underscore the vast potential of Diffusion Models in the field of video generation.
Beyond general video synthesis, the generation of human-centric videos has been the focal point of research, such as talking head. The objective of talking head is to generate the facial expressions from user-provided audio clips. Crafting these expressions involves capturing the subtleties and diversity of human facial movements, presenting a significant challenge in video synthesis. Traditional approaches often impose constraints on the final video output to simplify this task. For instance, some methods use 3D models to restrict facial keypoints, while others extract sequences of head movements from base videos to guide the overall motion. While these constraints reduce the complexity of video generation, they also tend to limit the richness and naturalness of the resulting facial expressions.
In this paper, our goal is to establish an innovative talking head framework designed to capture a broad spectrum of realistic facial expressions, including nuanced micro-expressions, and to facilitate natural head movements, thereby imbuing generated head videos with an unparalleled level of expressiveness. To achieve this goal, we propose a method that leverages the generative power of Diffusion Models, capable of directly synthesizing character head videos from a given image and an audio clip. This approach eliminates the need for intermediate representations or complex pre-processing, streamlining the creation of talking head videos that exhibit a high degree of visual and emotional fidelity, closely aligned with the nuances present in the audio input. Audio signals are rich in information related to facial expressions, theoretically enabling models to generate a diverse array of expressive facial movements. However, integrating audio with Diffusion Models is not a straightforward task due to the ambiguity inherent in the mapping between audio and facial expression. This issue can lead to instability in the videos produced by the model, manifesting as facial distortions or jittering between video frames, and in severe cases, may even result in the complete collapse of the video. To address this challenge, we have incorporated stable control mechanisms into our model, namely a speed controller and a face region controller, to enhance stability during the generation process. These two controllers function as hyperparameters, acting as subtle control signals that do not compromise the diversity and expressiveness of the final generated videos. Furthermore, to ensure that the character in the generated video remains consistent with the input reference image, we adopted and enhanced the approach of ReferenceNet by designing a similar module, FrameEncoding, aimed at preserving the character’s identity across the video.
@@
Finally, To train our model, we constructed a vast and diverse audio-video dataset, amassing over 250 hours of footage and more than 150 million images. This expansive dataset encompasses a wide range of content, including speeches, film and television clips, and singing performances, and covers multiple languages such as Chinese and English. The rich variety of speaking and singing videos ensures that our training material captures a broad spectrum of human expressions and vocal styles, providing a solid foundation for the development of EMO. We conducted extensive experiments and comparisons on the HDTF dataset, where our approach surpassed current state-of-the-art (SOTA) methods, including DreamTalk, Wav2Lip, and SadTalker, across multiple metrics such as FID, SyncNet, F-SIM, and FVD. In addition to quantitative assessments, we also carried out a comprehensive user study and qualitative evaluations, which revealed that our method is capable of generating highly natural and expressive talking and even singing videos, achieving the best results to date.
@@ 2 Related Work
Audio-driven talking head generation Audio-driven talking head generation can be broadly catgorized into two approaches:video-based methods[30, 5, 18] and single-image [33, 28, 15]. video-based talking head generation allows for direct editing on an input video segment. For example, Wav2Lip[18] regenerates lip movements in a video based on audio, using a discriminator for audio-lip sync. Its limitation is relying on a base video, leading to fixed head movements and only generating mouth movements, which can limit realism. For single-image talking head generation, a reference photo is utilized to generate a video that mirrors the appearance of the photo. [28] proposes to generate the head motion and facial expressions independently by learning blendshapes and head poses. These are then used to create a 3D facial mesh, serving as an intermediate representation to guide the final video frame generation. Similarly, [33] employs a 3D Morphable Model (3DMM) as an intermediate representation for generating talking head video. A common issue with these methods is the limited representational capacity of the 3D mesh, which constrains the overall expressiveness and realism of the generated videos. Additionally, both methods are based on non-diffusion models, which further limits the performance of the generated results. [15] attempts to use diffusion models for talking head generation, but instead of applying directly to image frames, it employs them to generate coefficients for 3DMM. Compared to the previous two methods, Dreamtalk offers some improvement in the results, but it still falls short of achieving highly natural facial video generation.
@@ 3.2 Network Pipelines
The overview of our method is shown in Figure 2. Our Backbone Network get the multi-frame noise latent input, and try to denoise them to the consecutive video frames during each time step, the Backbone Network has the similar UNet structure configuration with the original SD 1.5. 1) Similar to previous work, to ensure the continuity between generated frames, the Backbone Network is embedded with temporal modules. 2) To maintain the ID consistency of the portrait in the generated frames, we deploy a UNet structure called ReferenceNet parallel to the Backbone, it input the reference image to get the reference features. 3) To drive the character speaking motion, a audio layers is utilized to encode the voice features. 4) To make the motion of talking character controllable and stable, we use the face locator and speed layers to provide weak conditions.
@@
The pronunciation and tone in the voice is the main driven sign to the motion of the generated character. The features extracted from the input audio sequence by the various blocks of the pretrained wav2vec[22] are concatenated to yield the audio representation embedding, , for the frame. However, the motion might be influenced by the future/past audio segments, for example, opening mouth and inhaling before speaking. To address that, we define voice features of each generated frame by concatenating the features of nearby frames: , is the number of additional features from one side. To inject the voice features into the generation procedure, we add audio-attention layers performing a cross attention mechanism between the latent code and after each ref-attention layers in the Backbone Network.
@@
The ReferenceNet possesses a structure identical to that of the Backbone Network and serves to extract detailed features from input images. Given that both the ReferenceNet and the Backbone Network originate from the same original SD 1.5 UNet architecture, the feature maps generated by these two structures at certain layers are likely to exhibit similarities. Consequently, this facilitates the Backbone Network’s integration of features extracted by the ReferenceNet. Prior research[35, 9] has underscored the profound influence of utilizing analogous structures in maintaining the consistency of the target object’s identity. In our study, both the ReferenceNet and the Backbone Network inherit weights from the original SD UNet. The image of the target character is inputted into the ReferenceNet to extract the reference feature maps outputs from the self-attention layers. During the Backbone denoising procedure, the features of corresponding layers undergo a reference-attention layers with the extracted feature maps. As the ReferenceNet is primarily designed to handle individual images, it lacks the temporal layers found in the Backbone.
@@
Most works try inserting temporal mixing layers into the pretrained text-to-image architecture, to facilitate the understanding and encoding of temporal relationships between consecutive video frames. By doing so, the enhanced model is able to maintain continuity and consistency across frames, resulting in the generation of smooth and coherent video streams. Informed by the architectural concepts of AnimateDiff, we apply self-attention temporal layers to the features within frames. Specifically, we reconfigure the input feature map to the shape . Here, stands for the batch size, and indicate the spatial dimensions of the feature map, is the count of generated frames, and is the feature dimension. Notably, we direct the self-attention across the temporal dimension , to effectively capture the dynamic content of the video. The temporal layers are inserted at each resolution stratum of the Backbone Network. Most current diffusion-based video generation models are inherently limited by their design to produce a predetermined number of frames, thereby constraining the creation of extended video sequences. This limitation is particularly impactful in applications of talking head videos, where a sufficient duration is essential for the articulation of meaningful speaking. Some methodologies employ a frame from the end of the preceding clip as the initial frame of the subsequent generation, aiming to maintain a seamless transition across concatenated segments. Inspired by that, our approach incorporates the last frames, termed ’motion frames’ from the previously generated clip to enhance cross-clip consistency. Specifically, these motion frames are fed into the ReferenceNet to pre-extract multi-resolution motion feature maps. During the denoising process within the Backbone Network, we merge the temporal layer inputs with the pre-extracted motion features of matching resolution along the frames dimension. This straightforward method effectively ensures coherence among various clips. For the generation of the first video clip, we initialize the motion frames as zero maps.
@@
Specifically, as shown in Figure 2, we utilize a mask as the face region, which encompasses the facial bounding box regions of the video clip. We employ the Face Locator, which consists of lightweight convolutional layers designed to encode the bounding box mask. The resulting encoded mask is then added to the noisy latent representation before being fed into the Backbone. In this way, we can use the mask to control where the character face should be generated.
However, creating consistent and smooth motion between clips is challenging due to variations in head motion frequency during separate generation processes. To address this issue, we incorporate the target head motion speed into the generation. More precisely, we consider the head rotation velocity in frame and divide the range of speeds into discrete speed buckets, each representing a different velocity level. Each bucket has a central value and a radius . We retarget to a vector , where . Similar to the method used in the audio layers, the head rotation speed embedding for each frame is given by , and is subsequently processed by an MLP to extract speed features. Within the temporal layers, we repeat to the shape and implement a cross-attention mechanism that operates between the speed features and the reshaped feature map across the temporal dimension . By doing so and specifying a target speed, we can synchronize the rotation speed and frequency of the generated character’s head across different clips. Combined with the facial position control provided by the Face Locator, the resulting output can be both stable and controllable.
@@ 4.1 Implementations
We collected approximately 250 hours of talking head videos from the internet and supplemented this with the HDTF[34] and VFHQ[31] datasets to train our models. As the VFHQ dataset lacks audio, it is only used in the first training stage. We apply the MediaPipe face detection framework[13] to obtain the facial bounding box regions. Head rotation velocity was labeled by extracting the 6-DoF head pose for each frame using facial landmarks, followed by calculating the rotational degrees between successive frames.
The video clips sampled from the dataset are resized and cropped to . In the first training stage, the reference image and the target frame are sampled from the video clip separately, we trained the Backbone Network and the ReferneceNet with a batch size of 48. In the second and the third stage, we set as the generated video length, and the motion frames number is set to , we adopt a bath size of 4 for training. The additional features number is set to 2, following the Diffused Heads[27]. The learning rate for all stages are set to 1e-5. During the inference, we use the sampling algorithm of DDIM to generate the video clip for 40 steps, we assign a constant speed value for each frame generation. The time consumption of our method is about 15 seconds for one batch ( frames).
@@ 4.2 Experiments Setup
For methods comparisons, we partitioned the HDTF dataset, allocating 10% as the test set and reserving the remaining 90% for training. We took precautions to ensure that there was no overlap of character IDs between these two subsets.
We compare our methods with some previous works including: Wav2Lip[18], SadTalker[33], DreamTalk[15]. Additionally, we generated results using the released code from Diffused Heads[27], however, the model is trained on CREMA[1] dataset which contains only green background, the generated results are suboptimal. Furthermore, the results were compromised by error accumulation across the generated frames. Therefore, we only conduct qualitative comparison with the Diffused Heads approach. For DreamTalk, we utilize the talking style parameters as prescribed by the original authors.
To demonstrate the superiority of the proposed method, we evaluate the model with several quantitative metrics. We utilize Fréchet Inception Distance (FID)[7] to assess the quality of the generated frame[32]. Additionally, to gauge the preservation of identity in our results, we computed the facial similarity (F-SIM) by extracting and comparing facial features between the generated frames and the reference image. It is important to note that using a single, unvarying reference image could result in deceptively perfect F-SIM scores. Certain methods[18] might produce only the mouth regions, leaving the rest of the frame identical to the reference image, which could skew results. Therefore, we treat F-SIM as population-reference indices[27], with closer approximations to the corresponding ground truth (GT) values indicating better performance. We further employed the Fréchet Video Distance (FVD)[29] for the video-level evaluation. The SyncNet[2] score was used to assess the lip synchronization quality, a critical aspect for talking head applications. To evaluate the expressiveness of the facial expressions in the generated videos, we introduce the use of the Expression-FID (E-FID) metric. This involves extracting expression parameters via face reconstruction techniques, as described in [3]. Subsequently, we compute the FID of these expression parameters to quantitatively measure the divergence between the expressions in the synthesized videos and those in the ground truth dataset.

### v2

@@ Abstract
In this work, we tackle the challenge of enhancing the realism and expressiveness in talking head video generation by focusing on the dynamic and nuanced relationship between audio cues and facial movements. We identify the limitations of traditional techniques that often fail to capture the full spectrum of human expressions and the uniqueness of individual facial styles. To address these issues, we propose EMO, a novel framework that utilizes a direct audio-to-video synthesis approach, bypassing the need for intermediate 3D models or facial landmarks. Our method ensures seamless frame transitions and consistent identity preservation throughout the video, resulting in highly expressive and lifelike animations. Experimental results demonstrate that EMO is able to produce not only convincing speaking videos but also singing videos in various styles, significantly outperforming existing state-of-the-art methodologies in terms of expressiveness and realism.
@@ 1 Introduction
Diffusion Models have revolutionized the landscape of generative models, showcasing unparalleled capabilities in generating high-fidelity images[10, 25, 31, 5, 21]. These advancements have extended into video generation, sparking interest in leveraging these models for dynamic and engaging visual storytelling[11, 1, 39]. Beyond the general video generation, human-centric video generation, including portrait video generation[34, 33, 40, 19] and human animation[8, 12], has attracted considerable attention for its utility in creating digital human avatars and enhancing film production. A notably prominent area within this field is "talking head" video generation, which aims to synthesize a head video that matches with the input audio, while capturing the facial expressions, head movements and lip motion.
However, translating audio into head animation, such as facial expressions or head movements, is challenging due to the ambiguous and one-to-many mapping relationship. Most research on talking heads divides the process into head motion and facial expression components[34]. For head motion, some talking head techniques struggle to manage this aspect effectively and often resort to using a predefined pose sequence from an existing video[23, 20, 3] or use separate networks to individually address head poses and facial expressions[34]. In terms of facial expressions, some approaches opt for explicit intermediate signals like 3D face models or 2D facial landmarks to guide the generation[34, 40]. While these methods enhance the fidelity of specific aspects like lip synchronization, they tend to restrict the overall expressiveness and naturalness of the generated content. For example, the subtleties of spontaneous gestures or the nuanced expressions linked to speech’s emotional tone are often lost in translation, leading to less lifelike results. Therefore, to create highly expressive videos of talking head, it’s crucial to move away from the constraints of strong prior information and fully leverage the generative potential of the model.
In this paper, we introduce EMO (Emote Portrait Alive), a novel talking head framework that transforms a single portrait image into an expressive video, synchronized with audio, without relying on intermediate 3D representations or predefined motion templates. EMO harnesses the generative power of Diffusion Models to directly capture the intricate audio-visual correlations, facilitating the generation of dynamic and lifelike talking head videos. Specifically, our method extends Stable Diffusion for video by incorporating temporal modules and 3D convolutions, see Sec. 3.2.1. In order to lear the correlation between audio and videos, we introduce an audio feature extractor and employ attention module to modulate audio features into the backbone, see Sec. 3.2.2. To ensure stability without compromising expressiveness, we introduce novel mechanisms like the Face Locator and Speed Layers, to perform as week conditions for guiding the general area of the target face and the approximate speed level of movements. Unlike the strong priors, such as 3D intermediate representation, used by previous works, these conditions are not able to strictly constrains the face position and the speed of the generated videos, and therefore do not diminish the expressiveness of the generated videos (for more details, see Sec. 4.5). Additionally, we introduced Reference Net[12] to ensure consistent facial identity throughout the video (see Sec. 3.2.3) and implemented Motion Frame module to maintain continuity between adjacent video clips, enabling the generation of seamlessly infinite videos, Sec. 3.2.4.
@@
To train our model, we constructed a vast and diverse audio-video dataset, amassing over 250 hours of footage. This expansive dataset encompasses a wide range of content, including speeches, film and television clips, and singing performances, and covers multiple languages such as Chinese and English. The rich variety of speaking and singing videos ensures that our training material captures a broad spectrum of human expressions and vocal styles, providing a solid foundation for the development of EMO. We conducted extensive experiments across multiple datasets, comparing our model with various state-of-the-art methods. Our model outperforms these on several metrics, including a new metric we introduced, E-FID (Expression-FID), designed to evaluate the expressiveness of generated videos, thereby demonstrating superior performance. In summary, we make the following contributions: 1) We propose a novel talking head framework based on fully generative model without any 3D intermediates or pose prior. 2) Our method outperforms other existing SOTA methods in terms of expressiveness and realism of the talking head video. 3) We use weak conditions to stabilize the generation process without incurring significant loss in performance.
@@ 2 Related Work
Audio-driven talking head generation Audio-driven talking head generation can be broadly catgorized into two approaches:video-based methods[36, 6, 23, 29, 7] and single-image [40, 34, 19, 16]. video-based talking head generation allows for direct editing on an input video segment. For example, Wav2Lip[23] regenerates lip movements in a video based on audio, using a discriminator for audio-lip sync. Its limitation is relying on a base video, leading to fixed head movements and only generating mouth movements, which can limit realism. For single-image talking head generation, a reference photo is utilized to generate a video that mirrors the appearance of the photo. [34] proposes to generate the head motion and facial expressions independently by learning blendshapes and head poses. These are then used to create a 3D facial mesh, serving as an intermediate representation to guide the final video frame generation. Similarly, [40] employs a 3D Morphable Model (3DMM) as an intermediate representation for generating talking head video. A common issue with these methods is the limited representational capacity of the 3D mesh, which constrains the overall expressiveness and realism of the generated videos. Additionally, both methods are based on non-diffusion models, which further limits the performance of the generated results. [19] attempts to use diffusion models for talking head generation, but instead of applying directly to image frames, it employs them to generate coefficients for 3DMM. Compared to the previous two methods, Dreamtalk offers some improvement in the results, but it still falls short of achieving highly natural facial video generation.
@@ 3.2 Network Pipelines
The overview of our method is shown in Figure 2. Our Backbone Network get the multi-frame noise latent input, and try to denoise them to the consecutive video frames during each time step, the Backbone Network has the similar UNet structure configuration with the SD 1.5. 1) Similar to previous work, to ensure the continuity between generated frames, the Backbone Network is embedded with temporal modules. 2) To maintain the ID consistency of the portrait in the generated frames, we deploy a UNet structure called ReferenceNet parallel to the Backbone, it input the reference image to get the features. 3) To drive the character speaking motion, audio layers are utilized to encode the voice features. 4) To make the motion of talking character controllable and stable, we use the face locator and speed layers to provide weak conditions.
@@
The pronunciation and tone in the voice is the main driven sign to the generated character. The features extracted from the input audio sequence by the various blocks of the pretrained wav2vec[27] are concatenated to yield the audio representation embedding, , for the th frame. However, the motion might be influenced by the future/past audio segments, for example, opening mouth and inhaling before speaking. To address that, we define voice features of each generated frame by concatenating the features of nearby frames: , is the number of additional features from one side. To inject the voice features into the generation procedure, we add audio-attention layers performing a cross attention mechanism between the latent code and after each ref-attention layers in the Backbone Network.
@@
The ReferenceNet possesses a structure identical to that of the Backbone Network and serves to extract features from input images. Prior research[44, 12] has underscored the profound influence of utilizing analogous structures in maintaining the consistency of the target object’s identity. In our study, both the ReferenceNet and the Backbone Network inherit weights from the original SD UNet. The reference image is inputted into the ReferenceNet to extract the reference feature maps outputs from the self-attention layers. During the Backbone denoising procedure, the features of corresponding layers undergo a reference-attention layers with the extracted feature maps.
@@
Informed by the architectural concepts of AnimateDiff, we apply self-attention temporal layers to the features within frames. Specifically, we reconfigure the input feature map to the shape . Here, stands for the batch size, and indicate the spatial dimensions of the feature map, is the count of generated frames, and is the feature dimension. Notably, we direct the self-attention across the temporal dimension , to effectively capture the dynamic content of the video. The temporal layers are inserted at each resolution stratum of the Backbone Network. Most diffusion-based video generation models are inherently limited by their design to produce a predetermined number of frames, thereby constraining the creation of extended video. This limitation is particularly impactful in applications of talking head videos, where a sufficient duration is essential for the articulation of meaningful speaking. Some methodologies employ a frame from the end of the preceding clip as the initial frame of the subsequent generation, aiming to maintain a seamless transition across concatenated segments. Inspired by that, our approach incorporates the last frames, termed ’motion frames’ from the previously generated clip to enhance cross-clip consistency. Specifically, these motion frames are fed into the ReferenceNet to pre-extract multi-resolution motion feature maps. During the denoising process within the Backbone Network, we merge the temporal layer inputs with the pre-extracted motion features of matching resolution along the frames dimension. This straightforward method effectively ensures coherence among various clips. For the generation of the first video clip, we initialize the motion frames as zero maps.
@@
Specifically, as shown in Figure 2, we utilize a mask as the face region, which encompasses the facial bounding box (bbox) regions of the video clip. We employ the Face Locator, which consists of lightweight convolutional layers designed to encode the bounding box mask. The resulting encoded mask is then added to the noisy latent representation before being fed into the Backbone. We can use the mask to control where the character face should be generated.
However, creating consistent and smooth motion between clips is challenging due to variations in head motion frequency during separate generation processes. To address this issue, we incorporate the target head motion speed into the generation. More precisely, we consider the head rotation velocity in frame and divide the range of speeds into discrete speed buckets, each representing a different velocity level. Each bucket has a central value and a radius . We retarget to a vector , where the th value notated by . Similar to the method used in the audio layers, the head rotation speed embedding for each frame is given by . The speed embedding of each clip denoted by is subsequently processed by an MLP into a speed feature map . Within the temporal layers, we repeat to the shape and implement a cross-attention mechanism that operates between the speed features and the reshaped feature map across the temporal dimension . By doing so and specifying a target speed, we can synchronize the rotation speed and frequency of the generated character’s head across different clips. Combined with the facial position control provided by the Face Locator, the resulting output can be both stable and controllable.
@@ 4.1 Implementations
We collected approximately 250 hours of talking head videos from the internet and supplemented this with the HDTF[41] and VFHQ[37] datasets to train our models. As the VFHQ dataset lacks audio, it is only used in the first training stage. We apply the MediaPipe[17] to obtain the facial bbox. Head rotation velocity was labeled by extracting the 6-DoF head pose for each frame using facial landmarks, followed by calculating the rotational degrees between frames.
The video clips sampled from the dataset are resized and cropped to . In the first training stage, the reference image and the target frame are sampled from the video clip separately, we trained the Backbone Network and the ReferneceNet with a batch size of 48. In the second and the third stage, we set as the generated video length, and the motion frames number is set to , we adopt a bath size of 4 for training. The additional features number is set to 2. The learning rate for all stages are set to 1e-5. During the inference, we use the sampling algorithm of DDIM to generate the video clip for 40 steps, we assign a constant speed value for each frame generation. The time consumption of our method is about 15 seconds for one batch ( frames).
@@ 4.2 Experiments Setup
For methods comparisons, we partitioned the HDTF dataset, allocating 10% as the test set and reserving the remaining 90% for training. We took precautions to ensure that there was no overlap of character IDs between these two subsets. Additionally, to evaluate the methods in more variable scenarios, 1k video clips were extracted from the our collected internet video dataset, each with a duration of approximately 4 seconds. These clips predominantly feature expressive portrait videos with a substantial proportion depicting singing activities. Compared to the HDTF, the video dataset exhibits a broader diversity in terms of facial expressions, and the range of head motions.
We compare our methods with some previous works including: Wav2Lip[23], SadTalker[40], DreamTalk[19], MakeItTalk[42]. Additionally, we generated results using the released code from Diffused Heads[33], however, the model is trained on CREMA[2] dataset which contains only green background, the generated results are suboptimal. Furthermore, the results were compromised by error accumulation across the generated frames. Therefore, we only conduct qualitative comparison with the Diffused Heads approach. For DreamTalk, we utilize the talking style parameters as prescribed by the original authors.
To demonstrate the superiority of the proposed method, we evaluate the model with several metrics. We utilize Fréchet Inception Distance (FID)[9] to assess the quality of the generated frame[38]. Additionally, to gauge the preservation of identity in our results, we computed the facial similarity (F-SIM) by extracting and comparing facial features between the generated frames and the reference image. It is important to note that using a single, unvarying reference image could result in deceptively perfect F-SIM scores. Certain methods[23] might produce only the mouth regions, leaving the rest of the frame identical to the reference image, which could skew results. Therefore, we treat F-SIM as population-reference indices[33], with closer approximations to the corresponding ground truth (GT) values indicating better performance. We further employed the Fréchet Video Distance (FVD)[35] for the video-level evaluation. The SyncNet[3] score was used to assess the lip synchronization quality, a critical aspect for talking head applications. To evaluate the expressiveness of the facial expressions in the generated videos, we introduce the use of the Expression-FID (E-FID) metric. This involves extracting expression parameters via face reconstruction techniques, as described in [4]. Subsequently, we compute the FID of these expression parameters to quantitatively measure the divergence between the expressions in the synthesized videos and those in the ground truth dataset.